A service hatch in a drink water is positioned 1 m below the surface. The hatch dimensions are 700mm(w) x 800mm(h). The oke vertical side has two hinges, one at the top and one at the bottom. One fastening bolt is situated on the opposite sode 400mm from the top. Find the force in each of the hinges and the bolt.

HINT: First find the resultant force and where it acts, then use moments to find the various forces.


1
Expert's answer
2021-06-07T05:49:47-0400

"F= \rho g A h=1000*9.81*0.7*0.8*1= 5493.6 N"

Using moment, "M= F_{top} * \frac{h}{3}-F_{bottom} * \frac{h}{3}"

"M= 51.5* \frac{1.92}{3}-26.6 * \frac{1.48}{3}=19.99 N"